# Ambitions for theory in the physics of life

### v1

@@ 3 Matching distributions
In the fly embryo information flows from maternal inputs to the gap gene network to the striped patterns of pair rule gene expression. The previous lecture was about optimization at the output of the gap gene network. Can we also optimize the inputs? This is part of a more general question: given some signal processing system with fixed signal and noise characteristics, how can we choose the inputs to optimize information flow?
@@
The intuitions of Fig 16 can be formalized in the language of information theory. Information theory is a beautiful subject that has deep connections to statistical mechanics, but physicists vary in their level of comfort and familiarity with the ideas. I gave the lecture in Les Houches as if people knew the basics, but there was a strong desire for a more pedagogical introduction. So we followed the regular lecture with a tutorial, which we plan to have as an Appendix to the Summer School proceedings. [^25] My sincere thanks to Tarek Thome, who will co-author this Appendix, having recorded the tutorial and helped turn it into coherent prose. He even captured many of the excellent questions from the students. For completeness let me also note that (in contrast to many other subjects) one really can learn much of what you need to know about information theory by reading Shannon’s original paper [146]. The standard textbook is by Cover and Thomas [147], a version aimed at the physics community is by Mézard and Montanari [148], and I have tried to give a fuller account of these ideas in the context of biophysics [57]. Each of these texts provide more than you need to make sense out of these lectures, but perhaps you will want to explore more deeply.
@@ 3.1 One input, one output
The idea of using input in inverse proportion to their noise level is a precise version of familiar ideas. When writing we avoid words that we don’t know how to spell, and when speaking a foreign language we avoid constructions which we suspect we might get wrong. Different species of frogs call to one another in different frequency bands, and these match the bands where they hear best. In the low noise limit we can think of the matching condition as applying equally well to the input or the output,
@@ 3.2 Neural input/output relations
We can go one step further and choose such that $10<y<$ 1, so the optimized uniform distribution of outputs is $P_{Y}(y)=1$. In the low noise limit we have Eq (52), so that
@@
Two interesting things just happened. First, we started by optimizing the distribution of inputs and end up by expressing the result as an optimal input/output relation. Second, we have a prediction that the input/output relation should match the distribution of inputs, quantitatively and with no free parameters. If the distribution $P_{X}(x)$ has a single peak, $\bar{y}(x)$ should be roughly sigmoidal, as one finds for the input/output behaviors of many biological systems.
@@
Figure 18: Input/output relations match the distribution of inputs [149]. Brief changes in light intensity $\Delta I$ above or below background $\bar{I}$ produce transient voltage changes in the large monopolar cells (inset), and the peaks of these responses are taken as the cell’s output $\Delta V$ Normalized responses are compared to the cumulative probability distribution of light intensities, testing the predictions of Eq (65).
@@
There are many over–simplifications here. But the idea is powerful. When we ask why the input/output relation of a neuron looks the way it does, the standard answer is to start explaining all the molecular, cellular, and synaptic mechanisms that lead to the final phenomenology. Such explanations, as emphasized in the introductory lecture, lead us into a forest of parameters. Worse yet, these parameters are adjustable, for example if the cell in question had expressed different ion channels, different neurotransmitter receptors, etc. The idea here is different: input/output relations have the form that they do because they are matched to their inputs, optimizing information transmission. The fact that the predictions of a such a potentially general, parameter free theory are even approximately right is very encouraging.
@@
where $\bar{r}$ is the mean probability per unit time or rate of generating spikes, and $g(\tau)$ is the linear response function. This is too restrictive, and easily runs up agains the constraint that $P_{\rm spike}\geq 0$. A natural extension is to say that only the filtered inputs are important, but these could be processed nonlinearly,
@@
Figure 19: Adaptive rescaling in the responses of a motion–sensitive neuron in the fly visual system [155]. Normalized neural response $F(s_{1},s_{2})$ from Eq (69) projected onto stimulus features $s_{1}$ in (A, B) and onto $s_{2}$ in (C, D). (A, C) Input/output functions are different when inputs are chosen from different distributions and plotted in physical units , as indicated by the different symbols; inset in (A) indicates the standard deviation of the input velocities in the four distributions. (B, D) Input/output relations collapse in normalized units, confirming the prediction of Eq (72).
Figures 19A and C show the input/output relation in physical units. We see that when the same velocity is chosen from a different distribution, the response of the neuron can differ by orders of magnitude. But if we rescale the inputs by their standard deviations, these enormous variations collapse onto a single function, as in Figs 19B and D. As an aside, we note that this approach maps the input/output relations over a dynamic range of $\sim 10^{3}$, and the rescaling is essentially perfect across this full range. Although harder to visualize, we can seem the same rescaling in the $(s_{1},s_{2})$ plane. This is exactly what we expect from Eq (72).
@@ 3.3 Positional error in the embryo
If the distribution of positions $P(x)$ is smooth, then the conditional distribution $P(x|\{g_{\rm i}\})$ becomes a Gaussian with a mean that is slightly shifted from $x_{\rm true}$ by terms related to the noise $\delta g$ and a variance $\sigma_{x}^{2}=1/B$. All of these approximations are self–consistent if $\sigma_{x}$ that we compute in this was comes out to be small enough.
It is useful to think of these results as a generalization of error propagation, as in the discussion surrounding Fig 17. If the position $x$ is encoded by a single variable $g$ that has variance $\sigma_{g}^{2}$, then the effective variance in $x$ is
@@
but this reduces to Eq (83) if we rotate into the eigenbasis of the covariance matrix $\hat{C}(x)$. Notice that in this small noise approximation, the effective positional noise $\sigma_{x}$ depends on the true position but not on the particular concentrations $\{g_{\rm i}\}$ that we happen to observe. Also, in contrast to the discussion of decoding in §2.3, this analysis is local, and does not address ambiguities; again, it makes sense in a small noise limit.
Figure 20: Effective positional noise levels from gap gene expression levels [120]. Calculations of $\sigma_{x}$ are done in the small noise limit, from Eq (84), using the data illustrated in Fig 11. As indicated in the inset, different curves correspond to including just $g_{1}$, the combinations $\{g_{1},\,g_{2}\}$ and $\{g_{1},\,g_{2},\,g_{3}\}$ and finally all four gap genes $\{g_{1},\,g_{2},\,g_{3},\,g_{4}\}$. With all the genes we find that the positional error is small along the entire anterior–posterior axis, and close to $\sigma_{x}/L=0.01$ (dashed line).
Everything that we need to compute the effective positional error is contained the data of Fig 11, where we see the mean expression levels $\langle g_{\rm i}(x)\rangle$ and covariance matrices $\hat{C}(x)$ as functions of position. Although Eq (84) includes multiple genes, we can start with just one, corresponding to $g_{1}$ in Fig 11. We see that with this one gene the positional noise drops to $\sigma_{x}/L\sim 0.01$ in a window near the middle of the embryo, but is much larger outside this window. As we add more of the gap genes, the very largest values of $\sigma_{x}$ are reduced dramatically, and we end up with a nearly uniform $\sigma_{x}/L\sim 0.01$ across the full length of the embryo. Among other things, this is small enough that our approximations above are self–consistent. One should also see this as quantifying the progressive improvement of the decoding maps as we add more genes in Fig 13.303030Although the analysis of positional errors came before the full decoding maps, by roughly five years [10, 120].
There are at least two distinct points of interest in Fig 20. First, the scale of the positional errors, $\sigma_{x}/L\sim 0.01$. This is the same as the scale of positional errors in pair rule stripe positions, and also in the placement of the “cephalic furrow” [166], which is the first macroscopic structural change that happens after the embryo is separated into discrete cells. The conventional perspective on this system has been that information flow from maternal inputs to gap genes to pair rule genes entails a gradual refinement, with noisy inputs ultimately generating a precise and reproducible pattern [167, 168]. We see here that the precision visible at the pair rule genes already is present in the gap genes, so information must be transformed and preserved rather than refined.
@@
On the other hand, if we are free to optimize $P_{X}(x)$ over the range $0<x<L$ then form Eq (61) the maximum information possible given the measured $\sigma_{x}(x)$ becomes
@@
where $T$ is a Lagrange multiplier and we are explicit that the mapping $\{g_{\rm i}\}\rightarrow C$ is probabilistic; this is an example of the information bottleneck problem [171]. To extract all the available information, that is to have $I(C;x)\approx I(\{g_{\rm i}\};x)$, requires readout mechanisms with a capacity $I(C;\{g_{\rm i}\})$ of at least $8\,{\rm bits}$ [169]. Further, the most efficient mechanisms involve $C$ being sensitive to combinations of the different expression levels; if we have separate sensors $g_{\rm i}\rightarrow C_{\rm i}$ they need to have vastly more capacity. Finally, the embryo is in a regime where the trading of $I(C;x)$ vs $I(C;\{g_{\rm i}\})$ has a universal form [170]. If we take seriously the idea that $C$ is something like the occupancy of binding sites, or the collective states of the “enhancers” into which these sites are grouped, then these arguments about trading bits provide a path to predict the molecular mechanisms that instantiate the optimal decoding strategies of §2.3.
@@ 4 Network architecture
In the flow of positional information from the maternal inputs to the gap genes to the pair rule genes, we have seen evidence for optimization in the distribution of inputs (§3.3) and in the processing of the outputs (§2.3). We now have to ask if we can apply optimization principles to the network itself. Doing this involves returning to the challenge laid out in the first lecture: realistic descriptions of the gap gene network require 50+ parameters, and are in some obvious sense very complicated. To approach optimization of such a complex network it will be useful to break off smaller pieces of the problem and gain intuition.
@@ 4.1 Linear filtering in neural networks
again to keep things simple we imagine the number of inputs and outputs are the same so i=1, ${\rm i}=1,\,2,\,\cdots,\,N$ and similarly for $\rm j$. Let the noise $\zeta_{\rm i}$ be Gaussian and independent at every site $\rm i$. In this case optimizing the network means choosing the transformation matrix $\hat{M}$ so maximize the mutual information between inputs and outputs. If the variance of the noise is fixed, then one can always increase the mutual information by increasing the magnitude of $\hat{M}$, which seems like cheating. To have a well defined problem we should bound the magnitudes, so that the scale of the noise has meaning in relation to the inputs; a conventional choice is to constrain the summed variances of the outputs. This leaves us with the optimization problem
@@
A simple but important example of these ideas is color vision. Recall that in daylight our vision is based on three kinds of cones, each tuned to a different range of photon energies. We can think of these as sensitive approximately to red, green, and blue, though you should never say this in front of someone who actually studies color vision—“red” is a percept, not the label of a cell type; the convention is to describe the different cones as sensitive to long, medium, and short wavelengths. It is an old idea that the three cone signals are processed in well defined combinations corresponding roughly to the percepts of luminance, red vs green, and blue vs yellow. The history of these ideas is complicated, with all sorts of heroic figures shouting at one another. Thing were confusing in part because the “opponent process” theory seemed to involve four color axes (red, green, blue, and yellow) while the alternative “trichromatic” theory involved only three (which we now attach to the three types of cones). It seems to have been Schrödinger who sorted this out in 1925 [174]. We now know that the three cone signals indeed are grouped together in the three combinations already in the retina, and it is these combined signals that are transmitted to the brain [175].
In 1983, Buchsbaum and Gottschalk suggested that the transformation into luminance, red vs green, and blue vs yellow might be the solution to the optimization problem we are discussing here [176]. In order to see if this makes sense, we need to know the covariance matrix of the three cone signals. This depends strongly on the absorption spectra of the cone pigments, with large positive correlations arising simply because these spectra overlap. But the cone signal statistics also depend on the spectral composition of the images that the eye is looking at. To get at these, Ruderman, Cronin, and Chiao used a hyperspectral camera take pictures in natural environments—woodlands, forests and rainforests, and a mangrove swamp [177]. In effect this camera does spectroscopy in each pixel of a digital image, and then these spectrally resolved images can be projected onto the known sensitivities of the individual cone pigments to estimate the photon capture in each of the three cones looking at the same point. In keeping with the discussion of adaptation above (§3.2) the cone signals $x_{\rm i}$ were taken as the log of the number of photon counts. [^34] We use our cones in bright light, so there is no worry about zero counts.
@@
We see that $y_{1}$ is the sum of the signals in the three cones, and the weights are equal in the second decimal place or better. Similarly $y_{3}$ is the difference between the signals in long and medium length cones, with almost no contribution from the short wavelength cones; this is the “red minus green” opponent channel. Finally $y_{2}$ combines the long and medium cone signals and subtracts the short wavelength signal, corresponding to “yellow minus blue.” As with the luminance channel $y_{1}$, the combinations of cone signals in the two opponent channels are in integer ratios, with one percent accuracy. The signals along the different channels have standard deviations in the ratio $\sigma_{\ell}:\sigma_{by}:\sigma_{rg}\sim 1:0.2:0.02$. While we have a great appreciation for the subtleties of color, most of the variance in the images that we is in the luminance channel. [^35] In fact this might be a cautionary tale about dimensionality reduction. We can capture $\sim 95\%$ of the variance in what we see with just one dimension, corresponding to a completely greyscale world. Embarrassingly, I don’t know if the combination of cone signals into neural signals has been measured precisely enough to test these predictions of integer ratios.
@@
Given the statistical structure of natural images, optimization of information flow predicts that there will be zero gain for zero spatial frequency, $|\tilde{M}({\mathbf{k}}=0)|=0$. The transformation $M_{\rm ij}$ thus serves as roughly as a spatial differentiator, so that the output $y_{\rm i}$ is large when the pattern $\{x_{\rm j}\}$ includes a sharp edge, and is small when the $\{x_{\rm j}\}$ are nearly uniform. Qualitatively these predictions are correct for neurons in the retina. The output cells, which carry information from eye to brain, have long been known to have “center–surround” receptive fields [181, 182]: the response of these neurons is driven by the difference between the average light intensity in a small region and the average over a larger surrounding region, and in many cases the two regions have equal weight so that the response is differentiating. Barlow understood that this sort of spatial differencing, also called “lateral inhibition” [183], would remove the redundancy between signals in neighboring photoreceptors, enhancing the transmission of information by limited numbers of action potentials along limited numbers of neurons [172, 173]; the analysis here translates this qualitative observation into equations.
@@
Direct measurements of the transformation $M$ between neurons are not so easy. Our linear models make more sense if the cells have graded voltage responses, as in the first stages of vision. In the fly retina, for example, both the photoreceptor cells and the next cells in line (large monopolar cells, LMCs from §3.2) have a large regime in which the average voltage $V(t)$ responds linear to time variations in light intensity $I(t)$ around a background $I_{0}$,
@@
defining a transfer function $T(\tau)$; it is more convenient to think about the Fourier transform $\tilde{T}(\omega)$. These transfer functions can be measure in both the receptors and LMCs, with results in Fig 22A. Then we can take the ratio to estimate the filter across the synapse,
@@
Figure 22: Filtering at the first synapse in fly vision. (A) Linear response of voltage to temporal variations in light intensity, as in Eq (118), for photoreceptors and large monopolar cells (LMC). (B) Effective transfer function between photoreceptors and LMCs, from Eq (119). Redrawn from Ref [185].

### v2

@@ 3 Matching distributions
In the fly embryo information flows from maternal inputs to the gap gene network to the striped patterns of pair–rule gene expression. The previous lecture was about optimization at the output of the gap gene network. Can we also optimize the inputs? This is part of a more general question: given some signal processing system with fixed signal and noise characteristics, how can we choose the inputs to optimize information flow?
@@
The intuitions of Fig 16 can be formalized in the language of information theory. Information theory is a beautiful subject that has deep connections to statistical mechanics, but physicists vary in their level of comfort and familiarity with the ideas. I gave the lecture in Les Houches as if people knew the basics, but there was a strong desire for a more pedagogical introduction. So we followed the regular lecture with a tutorial [146], which we plan to have as an Appendix to the Summer School proceedings. [^25] My sincere thanks to Tarek Tohme, who will co-author this Appendix, having recorded the tutorial and helped turn it into coherent prose. He even captured many of the excellent questions from the students. For completeness let me also note that (in contrast to many other subjects) one really can learn much of what you need to know about information theory by reading Shannon’s original paper [147]. The standard textbook is by Cover and Thomas [148], a version aimed at the physics community is by Mézard and Montanari [149], and I have tried to give a fuller account of these ideas in the context of biophysics [57]. Each of these texts provide more than you need to make sense out of these lectures, but perhaps you will want to explore more deeply.
@@ 3.1 One input, one output
The result that inputs should be used in inverse proportion to their noise level is a precise version of familiar ideas. When writing we avoid words that we don’t know how to spell, and when speaking a foreign language we avoid constructions which we suspect we might get wrong. Different species of frogs call to one another in different frequency bands, and these match the bands where they hear best. In the low noise limit we can think of the matching condition as applying equally well to the input or the output,
@@ 3.2 Neural input/output relations
We can go one step further and choose units such that $10<y<$ 1, so the optimized uniform distribution of outputs is $P_{Y}(y)=1$. In the low noise limit we have Eq (52), so that
@@
Two interesting things just happened. First, we started by optimizing the distribution of inputs and ended up by expressing the result as an optimal input/output relation. Second, we have a prediction that the input/output relation should match the distribution of inputs, quantitatively and with no free parameters. If the distribution $P_{X}(x)$ has a single peak, $\bar{y}(x)$ should be roughly sigmoidal, as one finds for the input/output behaviors of many biological systems.
@@
Figure 18: Input/output relations match the distribution of inputs. Brief changes in light intensity $\Delta I$ above or below background $\bar{I}$ produce transient voltage changes in the large monopolar cells (inset), and the peaks of these responses are taken as the cell’s output $\Delta V$ Normalized responses are compared to the cumulative probability distribution of light intensities, testing the predictions of Eq (65). Data redrawn from Ref [150], under the Creative Commons Attribution NC-ND 3.0 License.
@@
There are many over–simplifications here, but the idea is powerful. When we ask why the input/output relation of a neuron looks the way it does, the standard answer is to start explaining all the molecular, cellular, and synaptic mechanisms that lead to the final phenomenology. Such explanations, as emphasized in the introductory lecture, lead us into a forest of parameters. Worse yet, these parameters are adjustable, for example if the cell in question had expressed different ion channels, different neurotransmitter receptors, etc. The idea here is different: input/output relations have the form that they do because they are matched to their inputs, optimizing information transmission. The fact that the predictions from such a potentially general, parameter free theory are even approximately right is very encouraging.
@@
where $\bar{r}$ is the mean probability per unit time or rate of generating spikes, and $g(\tau)$ is the linear response function. This is too restrictive, and easily runs up against the constraint that $P_{\rm spike}\geq 0$. A natural extension is to say that only the filtered inputs are important, but these could be processed nonlinearly,
@@
Figure 19: Adaptive rescaling in the responses of a motion–sensitive neuron in the fly visual system. Normalized neural response $F(s_{1},s_{2})$ from Eq (69) projected onto stimulus features $s_{1}$ in (A, B) and onto $s_{2}$ in (C, D). (A, C) Input/output functions are different when inputs are chosen from different distributions and plotted in physical units; inset in (A) indicates the standard deviation of the input velocities in the four distributions. (B, D) Input/output relations collapse in normalized units, confirming the prediction of Eq (72). Reproduced from Ref [156], with permission, and thanks to N Brenner and RR de Ruyter van Steveninck.
Figures 19A and C show the input/output relation in physical units. We see that when the same velocity is chosen from a different distribution, the response of the neuron can differ by orders of magnitude. But if we rescale the inputs by their standard deviations, these enormous variations collapse onto a single function, as in Figs 19B and D. As an aside, we note that this approach maps the input/output relations over a dynamic range of $\sim 10^{3}$, and the rescaling is essentially perfect across this full range. Although harder to visualize, we can see the same rescaling in the $(s_{1},s_{2})$ plane. This is exactly what we expect from Eq (72).
@@ 3.3 Positional error in the embryo
If the distribution of positions $P(x)$ is smooth, then the conditional distribution $P(x|\{g_{\rm i}\})$ becomes a Gaussian with a mean that is slightly shifted from $x_{\rm true}$ by terms related to the noise $\delta g$ and a variance $\sigma_{x}^{2}=1/B$. All of these approximations are self–consistent if the values of $\sigma_{x}$ that we compute in this way comes out to be small enough.
It is useful to think of these results as a generalization of error propagation, as in the discussion surrounding Fig 17. If the position $x$ is encoded by a single variable $g$ that has variance $\sigma_{g}^{2}$, then the effective variance in $x$ is determined by
@@
but this reduces to Eq (83) if we rotate into the eigenbasis of the covariance matrix $\hat{C}(x)$. Notice that in this small noise approximation, the effective positional noise $\sigma_{x}$ depends on the true position but not on the particular concentrations $\{g_{\rm i}\}$ that we happen to observe. Also, in contrast to the discussion of decoding in §2.3, this analysis is local, and does not address ambiguities; again, it makes sense in a small noise limit, where ambiguities are resolved.
Figure 20: Effective positional noise levels from gap gene expression levels. Calculations of $\sigma_{x}$ are done in the small noise limit, from Eq (84), using the data illustrated in Fig 11. As indicated in the inset, different curves correspond to including just $g_{1}$, the combinations $\{g_{1},\,g_{2}\}$ and $\{g_{1},\,g_{2},\,g_{3}\}$ and finally all four gap genes $\{g_{1},\,g_{2},\,g_{3},\,g_{4}\}$. With all the genes we find that the positional error is small along the entire anterior–posterior axis, and close to $\sigma_{x}/L=0.01$ (dashed line). Reproduced from Ref [120], with thanks to JO Dubuis, G Tkačik, EF Wieschaus, and T Gregor.
Everything that we need to compute the effective positional error is contained in the data of Fig 11, where we see the mean expression levels $\langle g_{\rm i}(x)\rangle$ and covariance matrices $\hat{C}(x)$ as functions of position. Although Eq (84) includes multiple genes, we can start with just one, corresponding to $g_{1}$ in Fig 11. We see that with this one gene the positional noise drops to $\sigma_{x}/L\sim 0.01$ in a window near the middle of the embryo, but is much larger outside this window. As we add more of the gap genes, the very largest values of $\sigma_{x}$ are reduced dramatically, and we end up with a nearly uniform $\sigma_{x}/L\sim 0.01$ across the full length of the embryo. Among other things, this is small enough that our approximations above are self–consistent. One should also see this as quantifying the progressive improvement of the decoding maps as we add more genes in Fig 13.303030Although the analysis of positional errors came before the full decoding maps, by roughly five years [10, 120].
There are at least two distinct points of interest in Fig 20. First, the scale of the positional errors, $\sigma_{x}/L\sim 0.01$. This is the same as the scale of positional errors in pair–rule stripe positions, and also in the placement of the “cephalic furrow” [167], which is the first macroscopic structural change that happens after the embryo is separated into discrete cells. The conventional perspective on this system has been that information flow from maternal inputs to gap genes to pair–rule genes entails a gradual refinement, with noisy inputs ultimately generating a precise and reproducible pattern [168, 169]. We see here that the precision visible at the pair–rule genes already is present in the gap genes, so information must be transformed and preserved rather than refined.
@@
On the other hand, if we are free to optimize $P_{X}(x)$ over the range $0<x<L$ then from Eq (61) the maximum information possible given the measured $\sigma_{x}(x)$ becomes
@@
where $T$ is a Lagrange multiplier and we make explicit that the mapping $\{g_{\rm i}\}\rightarrow C$ is probabilistic; this is an example of the information bottleneck problem [172]. To extract all the available information, that is to have $I(C;x)\approx I(\{g_{\rm i}\};x)$, requires readout mechanisms with a capacity $I(C;\{g_{\rm i}\})$ of at least $8\,{\rm bits}$ [170]. Further, the most efficient mechanisms involve $C$ being sensitive to combinations of the different expression levels; if we have separate sensors $g_{\rm i}\rightarrow C_{\rm i}$ they need to have vastly more capacity. Finally, the embryo is in a regime where the trading of $I(C;x)$ vs $I(C;\{g_{\rm i}\})$ has a universal form [171]. If we take seriously the idea that $C$ is something like the occupancy of binding sites, or the collective states of the “enhancers” into which these sites are grouped, then these arguments about trading bits provide a path to predict the molecular mechanisms that instantiate the optimal decoding strategies of §2.3.
@@ 4 Network architecture
In the flow of positional information from the maternal inputs to the gap genes to the pair–rule genes, we have seen evidence for optimization in the distribution of inputs (§3.3) and in the processing of the outputs (§2.3). We now have to ask if we can apply optimization principles to the network itself. Doing this involves returning to the challenge laid out in the first lecture: realistic descriptions of the gap gene network require 50+ parameters, and are in some obvious sense very complicated. To approach optimization of such a complex network it will be useful to break off smaller pieces of the problem and gain intuition.
@@ 4.1 Linear filtering in neural networks
again to keep things simple we imagine the number of inputs and outputs are the same so i=1, ${\rm i}=1,\,2,\,\cdots,\,N$ and similarly for $\rm j$. Let the noise $\zeta_{\rm i}$ be Gaussian and independent at every site $\rm i$. In this case optimizing the network means choosing the transformation matrix $\hat{M}$ to maximize the mutual information between inputs and outputs. If the variance of the noise is fixed, then one can always increase the mutual information by increasing the magnitude of $\hat{M}$, which seems like cheating. To have a well defined problem we should bound the magnitudes, so that the scale of the noise has meaning in relation to the inputs; a conventional choice is to constrain the summed variances of the outputs. This leaves us with the optimization problem
@@
A simple but important example of these ideas is color vision. Recall that in daylight our vision is based on three kinds of cones, each tuned to a different range of photon energies. We can think of these as sensitive approximately to red, green, and blue, though you should never say this in front of someone who actually studies color vision—“red” is a percept, not the label of a cell type; the convention is to describe the different cones as sensitive to long, medium, and short wavelengths. It is an old idea that the three cone signals are processed in well defined combinations corresponding roughly to the percepts of luminance, red vs green, and blue vs yellow. The history of these ideas is complicated, with all sorts of heroic figures shouting at one another. Things were confusing in part because the “opponent process” theory seemed to involve four color axes (red, green, blue, and yellow) while the alternative “trichromatic” theory involved only three (which we now attach to the three types of cones). It seems to have been Schrödinger who sorted this out in 1925 [175]. We now know that the three cone signals indeed are grouped together in the three combinations already in the retina, and it is these combined signals that are transmitted to the brain [176].
In 1983, Buchsbaum and Gottschalk suggested that the transformation into luminance, red vs green, and blue vs yellow might be the solution to the optimization problem we are discussing here [177]. In order to see if this makes sense, we need to know the covariance matrix of the three cone signals. This depends strongly on the absorption spectra of the cone pigments, with large positive correlations arising simply because these spectra overlap. But the cone signal statistics also depend on the spectral composition of the images that the eye is looking at. To get at these, Ruderman, Cronin, and Chiao used a hyperspectral camera to take pictures in natural environments—woodlands, forests and rainforests, and a mangrove swamp [178]. In effect this camera does spectroscopy in each pixel of a digital image, and then these spectrally resolved images can be projected onto the known sensitivities of the individual cone pigments to estimate the photon capture in each of the three cones looking at the same point. In keeping with the discussion of adaptation above (§3.2) the cone signals $x_{\rm i}$ were taken as the log of the number of photon counts. [^34] We use our cones in bright light, so there is no worry about zero counts.
@@
We see that $y_{1}$ is the sum of the signals in the three cones, and the weights are equal in the second decimal place or better. Similarly $y_{3}$ is the difference between the signals in long and medium length cones, with almost no contribution from the short wavelength cones; this is the “red minus green” opponent channel. Finally $y_{2}$ combines the long and medium cone signals and subtracts the short wavelength signal, corresponding to “yellow minus blue.” As with the luminance channel $y_{1}$, the combinations of cone signals in the two opponent channels are in integer ratios, with one percent accuracy. The signals along the different channels have standard deviations in the ratio $\sigma_{\ell}:\sigma_{by}:\sigma_{rg}\sim 1:0.2:0.02$. While we have a great appreciation for the subtleties of color, most of the variance in the images that we see is in the luminance channel. [^35] In fact this might be a cautionary tale about dimensionality reduction. We can capture $\sim 95\%$ of the variance in what we see with just one dimension, corresponding to a completely greyscale world. Embarrassingly, I don’t know if the combination of cone signals into neural signals has been measured precisely enough to test these predictions of integer ratios.
@@
Given the statistical structure of natural images, optimization of information flow predicts that there will be zero gain for zero spatial frequency, $|\tilde{M}({\mathbf{k}}=0)|=0$. The transformation $M_{\rm ij}$ thus serves roughly as a spatial differentiator, so that the output $y_{\rm i}$ is large when the pattern $\{x_{\rm j}\}$ includes a sharp edge, and is small when the $\{x_{\rm j}\}$ are nearly uniform. Qualitatively these predictions are correct for neurons in the retina. The output cells, which carry information from eye to brain, have long been known to have “center–surround” receptive fields [182, 183]: the response of these neurons is driven by the difference between the average light intensity in a small central region and the average over a larger surrounding region, and in many cases the two regions have equal weight so that the response is differentiating. Barlow understood that this sort of spatial differencing, also called “lateral inhibition” [184], would remove the redundancy between signals in neighboring photoreceptors, enhancing the transmission of information by limited numbers of action potentials along limited numbers of neurons [173, 174]; the analysis here translates this qualitative observation into equations.
@@
Direct measurements of the transformation $M$ between neurons are not so easy. Our linear models make more sense if the cells have graded voltage responses, as in the first stages of vision. In the fly retina, for example, both the photoreceptor cells and the next cells in line (large monopolar cells, LMCs from §3.2) have a large regime in which the average voltage $V(t)$ responds linearly to time variations in light intensity $I(t)$ around a background $I_{0}$,
@@
defining a transfer function $T(\tau)$; it is more convenient to think about the Fourier transform $\tilde{T}(\omega)$. These transfer functions can be measured in both the receptors and LMCs, with results in Fig 22A. Then we can take the ratio to estimate the filter across the synapse,
@@
Figure 22: Filtering at the first synapse in fly vision. (A) Linear response of voltage to temporal variations in light intensity, as in Eq (118), for photoreceptors and large monopolar cells (LMC). (B) Effective transfer function between photoreceptors and LMCs, from Eq (119). My thanks to SB Laughlin and RR de Ruyter van Steveninck for sharing these data, from experiments described in Ref [186].